4.   A perfectly competitive firm has total revenue and

total cost curves given by:








            TR = 100Q                           TC = 5,000 + 2Q +

0.2 Q2








a.         Find the profit-maximizing output for

this firm.




b.         What profit does the firm make?










1
Expert's answer
2022-01-10T09:55:37-0500

a.The equality of MR and MC is a condition for maximizing profits for any firm, regardless of the market structure in which it operates (perfect or imperfect competition).

"MR=TR'=(100Q)'=100"

"MC=TC'=(5000+2Q+0.2Q^2)'=2+0.4Q"

100=2+0.4Q

98=0.4Q

Q=245

P=MC

MC=2+0.4Q

MC=P=100


b.

"Profit=TR-TC=100\times245-(5000+2\times245+0.2(245)^2)=24500-17495=7005"